OpenStudy (anonymous):

Which cosine function has a maximum of 2, a minimum of -2, and a period of 2pi/3? y=3cos2(theta) y=2cos(2theta)/(3) y=2cos3(theta) y=2cos(theta)/(3)

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@johnweldon1993

hartnn (hartnn):

Amplitude = A = 2 is given Period T = 2pi/3 use the general form of cos function : \(\Large y = A \cos (\dfrac{2\pi}{T}\theta )\)

OpenStudy (anonymous):

So it would be 2 cos 2theta/3?

hartnn (hartnn):

how did you get 2theta/3 ?? 2pi/T with T =2pi/3 gives 2pi/(2pi/3) = ... ?

OpenStudy (anonymous):

I entered it into my calculator. But 2pi/2pi/3 is 3

hartnn (hartnn):

yes, 3 is correct!

hartnn (hartnn):

y=2cos3(theta)

OpenStudy (anonymous):

Ohh okay! Thank you! :) These are difficult for me. I've been practicing a little more.

hartnn (hartnn):

good! keep practicing :) welcome ^_^